OpenStudy (anonymous):

if cos0 (line in the middle) = -5/13 and sin 0(line in the middle) <0 , what is tan0 (line in the middle) a. 12/5 b. 13/5 c. -12/5 d. sqrt 194/5

OpenStudy (anonymous):

line in the middle as in \(\large \color{red}{\theta}\) ?

OpenStudy (anonymous):

as if trig wasn't annoying enough they have to use greek letters that one is called "theta" we can solve this by drawing a triangle with the adjacent side 5 and the hypotenuse 13 like this 

OpenStudy (anonymous):

all you need is the opposite side, which, by pythagoras, is \[\sqrt{13^2-5^2}=\sqrt{136-25}=\sqrt{144}=12\]

OpenStudy (anonymous):

typo there, should be \(\sqrt{169-25}=\sqrt{144}\)

OpenStudy (anonymous):

then the tangent is "opposite over adjacent" i.e. \(\frac{12}{5}\)

OpenStudy (anonymous):

thankyou and sorry i dont know how to do the sign for theta and kno one would tell me what theta was :p

OpenStudy (anonymous):

yw it is just a greek letter that they often use for angles, like alpha \(\alpha\) and beta \(\beta\) not sure why they use greek ones

OpenStudy (anonymous):

how do u type it out on the computer?

OpenStudy (anonymous):

i write the code in something called \(\LaTeX\) if you want to see the code, right click on what you want to view and select "show math as" then "tex commands" alpha and beta are easy, they are \alpha and \beta inclosed in \(

OpenStudy (kc_kennylau):

You can copy this and paste it wherever you want: θ

OpenStudy (kc_kennylau):

(I enabled the Greek keyboard)

OpenStudy (anonymous):

thankyou kc_kennylau

OpenStudy (anonymous):

i wouldn't worry about it too much, you might as well use x it makes no difference

OpenStudy (anonymous):

oh ok thankyou